OpenStudy (anonymous):

solve the following ordinary differential eq. x*(d^2u/dx^2) = 5*(du/dx) Hint: substitute v= du/dx

OpenStudy (anonymous):

help!

OpenStudy (dumbcow):

make the substitution, notice 2nd derivative of u equals 1st derivative of v \[x*\frac{dv}{dx} = 5v\] then separate variables \[\frac{dv}{v} = \frac{5}{x} dx\] then integrate both sides to solve for v

OpenStudy (dumbcow):

once you have v, integrate again to get u note: you will have 2 constants of integration in answer